OpenStudy (ttop0816):

The range of y = 2/3 sin x for π ≤ x ≤ 3pi/2 is _____. -2/3 ≤ y ≤ 2/3 -2/3 ≤ y ≤ 0 0 ≤ y ≤ 2/3 2/3 ≤ y ≤ 1

hartnn (hartnn):

whats the range of sin x ? you know ?

OpenStudy (ttop0816):

yes it is -1≤ y ≤1

hartnn (hartnn):

yes, so range of 2/3 sin x will be from -2/3≤ y ≤2/3 but there's an interval given. so, plug the endpoints in your equation, put x= pi in y = 2/3 sin x, find 'y1' put x= 3pi/2 in y = 2/3 sin x, find 'y2' and you will get the end-points of your range. [y1,y2]

OpenStudy (ttop0816):

how could i solve this in a graph form? or can i solve it in a graph form? because thats how im learning at school ):

hartnn (hartnn):

2/3 sin pi = ... ? 2/3 sin 3p/2 = ... ?

OpenStudy (ttop0816):

2/3 sin pi= 0?

OpenStudy (ttop0816):

tbh i dont get what you have to plug in here ): put x= pi in y = 2/3 sin x, find 'y1' put x= 3pi/2 in y = 2/3 sin x, find 'y2' and you will get the end-points of your range. [y1,y2] what is x here?

hartnn (hartnn):

you can graph cos x, right ? just make the amplitude = 2/3 and see what values it takes from pi to 3pi/2

OpenStudy (ttop0816):

0 ≤ y ≤ 2/3 maybe??

hartnn (hartnn):

when x= 3pi/2 , y = -2/3 so the range is actually -2/3 to 0

OpenStudy (ttop0816):

ohhhhhh! i got it hope i remember this when im solving other problems :P

hartnn (hartnn):

god luck! :)